Prove that a group has exactly one element with the property that


x * x = x (this is called an idempotent element)



1
Expert's answer
2021-02-24T06:52:14-0500

Let "x*x=x" for any element "x" of a group "G." Let "e" be an identity element of the group "G". Let "x\in G" be arbitrary. It follows that "e*x=x". On the other hand, "x*x=x". Therefore, "x*x=e*x." The right cancellation property of a group implies that "x=e" for all "x\in G". Consequently, "G" has exactly one element.